a store normally sells cashews for $9.00 per pound and almonds for $3.50 per pound. But at the end of the month the almonds had not sold well, so, in order to sell 60lb of almonds, the manager decided to mix the 60lb of almonds with some cashews and sell the mixture for $7.50 per pound. How many pounds of cashews should be mixed with almonds to ensure no change in the profit
Answer by stanbon(57225)

a store normally sells cashews for $9.00 per pound and almonds for $3.50 per pound.
But at the end of the month the almonds had not sold well, so, in order to sell 60lb of almonds, the manager decided to mix the 60lb of almonds with some cashews and sell the mixture for $7.50 per pound.
How many pounds of cashews should be mixed with almonds to ensure no change in the profit
---------------
Almond DATA:
Amt. = 60 lbs. ; Value = 3.5*60 = 210 dollars
---------------
Cashew DATA:
Amt. = x lbs. ; value = 9x dollars
--------------
Mixture DATA:
Amt. = 60+x lbs; value = 7.5*(60+x) = 450 + 7.5x dollars
------------------
EQUATION:
value + value = value
210 + 9x = 450+7.5x
1.5x = 240
x = 160 lbs. (amount of cashews that should be added to make the mixture)
==================
Cheers,
Stan H.